OpenStudy (anonymous):

Write a quadratic function that fits the given set of points. (-2, 5), (0, -3), and (3, 0)

OpenStudy (***[isuru]***):

hey... take the pair (0 , -3) so if the function is y = ax^2 + bx + c then -3 = a(0)^2 +b(0) + c which means c = -3 take the pair (3 , 0) y = ax^2 + bx + c 0 =a(3)^2 + 3b -3 9a + 3b = 3 -----(1) now... go to the pair (-2 , 5) y = ax^2 + bx + c 5 = a(-2)^2 -2b -3 4a - 2b = 8 -----(2) now multiply (1) by 2 and (2) by 3... ( 1) x 2 18a + 6b = 6 ----(3) (2) x 3 12a - 6b = 24 ----(4) (3) + (4) 30 a = 30 a = 1 substitute a = 1 in (2) 4 - 2b = 8 2b = -4 b = -2 so.... the function is y = x^2 -2x -3 hope this will help ya!!